OpenStudy (anonymous):

Express each quotient in the simplest form. 10xy^2 / 3z divide 5xy / 6z^3 a. 2yz^2 b. 2y^2z^2 c. 4yz^2 d. 4y^2z^2 help me please.

OpenStudy (dan815):

a/b divide by c/d = a/b* d/c <-- you can flip the division like that

OpenStudy (dan815):

can you do that with your expressions..

OpenStudy (jdoe0001):

same as before, then cancel out what you can and simplify keep in mind that \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\)

OpenStudy (anonymous):